# Consistency of Neural Causal Partial Identification

### v1

@@ B.1 Proof of Theorem 1
$\int g(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k})\ d\left(\mu_{V_{k}|\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}-\hat{\mu}_{\hat{V}_{k}|\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right)$
$=g\left(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k-1},f_{k}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)\right)-g\left(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k-1},\hat{f}_{k}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)\right)$
$\leqslant\left\|f_{k}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)-\hat{f}_{k}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)\right\|_{\infty}\leqslant S,$ (9)
where we have used the Lipschitz property of $g$ in the first inequality. If $V_{k}$ is categorical, let $\hat{p}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)=(\hat{p}_{1},\cdots,\hat{p}_{n_{i}})={\hat{f}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)}$ and $p\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)=(p_{1},\cdots,p_{n_{i}})=f\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)$ , we get
$\int g(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k})\ d\left(\mu_{V_{k}|\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}-\hat{\mu}_{\hat{V}_{k}|\widehat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right)$
$=\sum_{k=1}^{n_{i}-1}(g(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,k)\ -g(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,n_{i}))(p_{k}-\hat{p}_{k})\leqslant K\sum_{k=1}^{n_{i}-1}|p_{k}-\hat{p}_{k}|\leqslant K^{2}S,$ (10)
@@
$=\int g\left(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k-1},{f}_{k}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)\right)\ d\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}-\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right).$ (11)
Since $g,{f}_{k}$ are Lipschitz continuous functions, $g\left(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k-1},{f}_{k}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)\right)$ is $(L+1)$-Lipschitz continuous with respect to $(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k-1})$. We have
$\int g\left(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k-1},{f}_{k}\left(\text{pa}(v_{k}),\boldsymbol{u}_{V_{k}}\right)\right)$ $d\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}-\hat{\mu}_{\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right)$
$\leqslant(L+1)W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right).$ (12)
@@
$\quad=\sum_{k=1}^{n_{i}}{p}_{k}\int g(\boldsymbol{u},\boldsymbol{t},x_{1},\cdots,k)\ d\left(\mu_{\boldsymbol{U},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}-\hat{\mu}_{\hat{\boldsymbol{U}},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right).$
@@
Since $g,f_{k}$ are $L$-Lipschitz continuous functions, $g(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k-1},i)$ is $L$-Lipschitz continuous with respect to $(\boldsymbol{u},\boldsymbol{t},v_{1},\cdots,v_{k-1})$. We have
$\sum_{k=1}^{n_{i}}\int{p}_{k}g(\boldsymbol{u},\boldsymbol{t},x_{1},\cdots,k)\ d\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}-\hat{\mu}_{\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right)$ $\leqslant LW\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:k-1}^{\boldsymbol{T}}\right).$ (13)
Combine (9)-(13), we prove Equation 8. By induction, one can easily get
$W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:n_{V}-n_{T}}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:n_{V}-n_{T}}^{\boldsymbol{T}}\right)$ $\leqslant(L+1)^{n_{V}-n_{T}}\left(W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T}}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T}}^{\boldsymbol{T}}\right)+2K^{2}S/L\right)-2K^{2}S/L$
$=\frac{(L+1)^{n_{V}-n_{T}}-1}{L}2K^{2}S+(L+1)^{n_{V}-n_{T}}W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T}}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T}}^{\boldsymbol{T}}\right)$
@@
Let $C_{\mathcal{G}}(L,K)=2K^{2}\cdot\max\{\frac{(L+1)^{n_{V}-n_{T}}-1}{L},(L+1)^{n_{V}-n_{T}}\}$ and notice that $W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T}}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T}}^{\boldsymbol{T}}\right)=W\left(P^{\mathcal{M}}(\boldsymbol{U}),P^{\hat{\mathcal{M}}}(\hat{\boldsymbol{U}})\right)$ because interventions $\boldsymbol{T}=\boldsymbol{t}$ are the same, we get
$W\left(P^{\mathcal{M}}(\boldsymbol{V}(\boldsymbol{t})),P^{\hat{\mathcal{M}}}(\boldsymbol{V}(\boldsymbol{t}))\right)$ $\leqslant W\left(\mu_{\boldsymbol{U}_{0},\boldsymbol{T},1:n_{V}-n_{T}}^{\boldsymbol{T}},\hat{\mu}_{\hat{\boldsymbol{U}}_{0},\boldsymbol{T},1:n_{V}-n_{T}}^{\boldsymbol{T}}\right)$
$\leqslant C_{\mathcal{G}}(L,K)\left(S+W\left(P^{\mathcal{M}}(\boldsymbol{U}),P^{\hat{\mathcal{M}}}(\hat{\boldsymbol{U}})\right)\right).$ (14)
B.2 Proof of results in Section 3.1
Proof of Theorem 2.
@@
Recall that the neural network consists of three parts $\hat{g}=\hat{g}_{3}^{\tau}\circ\hat{g}_{2}\circ\hat{g}_{1}$ and $\hat{g}_{2},\hat{g}_{1}$ have a separable form, dealing with each coordinate of the input individually. As mentioned before, each coordinate approximates the distribution of one connected component of the support. We first construct $\hat{g}_{1}$ and $\hat{g}_{2}$, then use Gumbel-Softmax layer to combine each component together.
@@
To construct the first two parts, we only need to consider each coordinate individually. For the $i$-th component $C_{i}$, let $\mu_{i}=\mathbb{P}_{C_{i}}/\mathbb{P}(C_{i})$, where $\mathbb{P}_{C_{i}}$ is measure $\mathbb{P}$ restricted to component $C_{i}$. Under 4, there exists a Lipschitz map $g_{2}^{i}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$. By [42, Theorem VIII.1], there exists quantized ReLu network $\hat{g}_{1}^{i}$ of width $W_{1}$ and depth $\Theta(d_{i}^{C})$ (let $s=\frac{1}{\lceil n(n-1)\rceil}$ in [42, Theorem VIII.1]), such that
$W\left(\left((g_{2}^{i})^{-1}\right)_{\#}\mu_{i},P(\hat{g}_{1}^{i}(U_{i}))\right)\leqslant O\left(W_{1}^{-1/d_{i}^{C}}\right),$
@@
where $U_{i}$ are i.i.d. uniform random variables on $[0,1]$. By [52, Theorem 2], there exist a deep ReLu network $\hat{g}_{2}^{i,j}$ of width $\Theta\left(d_{i}^{C}\right)$ and depth $L_{2}$ such that
$\|\left(g_{2}^{i}\right)_{j}-\hat{g}_{2}^{i,j}\|_{\infty}\leqslant O\left(L_{2}^{-2/d_{i}^{C}}\right),$
Let $\hat{g}_{2}^{i}(x)=\left(\hat{g}_{2}^{i,1}(x),\cdots,\hat{g}_{2}^{i,d}(x)\right)$, we get
$\|g_{2}^{i}-\hat{g}_{2}^{i}\|_{\infty}\leqslant O\left(L_{2}^{-2/d_{i}^{C}}\right).$
Thus, the width of $\hat{g}_{2}^{i}$ is $\Theta\left(d\cdotp d_{i}^{C}\right)$. Let
$\hat{g}_{i}(x)=\left(\hat{g}_{i}^{1}(x_{1}),\cdots,\hat{g}_{i}^{N_{C}}(x_{N_{C}})\right),i=1,2,x_{k}\in\mathbb{R}^{d_{i}},d_{1}=1,d_{2}=d,k=1,\cdots,N_{C},$
where $N_{C}$ is the number of connected components. By Lemma 3, we get
$W\left(\mu_{i},P(\hat{g}_{2}^{i}\circ\hat{g}_{1}^{i}(U_{i}))\right)$ $=W\left(\left(g_{2}^{i}\right)_{\#}\left(g_{2}^{i}\right)^{-1}_{\#}\mu_{i},\left(\hat{g}_{2}^{i}\right)_{\#}P(\hat{g}_{1}^{i}(U_{i}))\right)$
$\leqslant LW\left(\left(g_{2}^{i}\right)^{-1}_{\#}\mu_{i},P(\hat{g}_{1}^{i}(U_{i}))\right)+\|g_{2}^{i}-\hat{g}_{2}^{i}\|_{\infty}$
@@
Next, let $p_{i}=\mathbb{P}(C_{i})$ and the distribution of $\hat{g}_{2}^{k}\circ\hat{g}_{1}^{k}(U_{k})$ be $\hat{\mu}_{k}$, then we have $\mathbb{P}=\sum_{k=1}^{N_{C}}p_{k}\mu_{k}$. By Lemma 4, we have
$W\left(\mathbb{P},\sum_{k=1}^{N_{C}}p_{k}\hat{\mu}_{k}\right)$ $\leqslant\sum_{k=1}^{N_{C}}p_{k}W\left(\mu_{i},\hat{\mu}_{k}\right)=O\left(W_{1}^{-1/\max\{d_{i}^{C}\}}+L_{2}^{-2/\max\{d_{i}^{C}\}}\right).$
@@
By Lemma 5,
$W\left(\left(\nu^{\tau},P(\left(\hat{g}_{2}^{i}\circ\hat{g}_{1}^{i}\right)(U_{i}))\right),\left(\nu^{0},P(\left(\hat{g}_{2}^{i}\circ\hat{g}_{1}^{i}\right)(U_{i}))\right)\right)\leqslant O\left(\tau-\tau\log\tau\right).$ (15)
@@
$\hat{g}_{3}^{\tau}(x_{1},\cdots,x_{N_{C}})=\sum_{k=1}^{N_{C}}w_{k}^{\tau}\cdotp x_{k},\quad x_{k}\in\mathbb{R}^{d},\ w^{\tau}=\left(w_{1}^{\tau},\cdots,w_{N_{C}}^{\tau}\right).$
We denote $g^{0}_{3}=\lim_{\tau\rightarrow 0}g^{\tau}_{3}$. By 4, the support is bounded. Thus, $g_{2}^{i}$ and $\hat{g}_{2}^{i}$ are bounded, i.e., $\|\hat{g}_{2}^{i}\ (x)\|_{\infty}\leqslant K$. Since functions $h(w,X)=w^{\mathsf{T}}X$ is Lipschitz continuous in the region $\|w\|_{1}=1,\|X\|_{\infty}\leqslant K$, we get
$W\left(P(\hat{g}_{3}^{0}\circ\hat{g}_{2}\circ\hat{g}_{1}(U_{k})),P(\hat{g}_{3}^{\tau}\circ\hat{g}_{2}\circ\hat{g}_{1}(U_{k}))\right)\leqslant O\left(\tau-\tau\log\tau\right)$ (16)
from (15). Putting things together, we get
$W\left(\mathbb{P},P(\hat{g}_{3}^{\tau}\circ\hat{g}_{2}\circ\hat{g}_{1}(U_{k}))\right)$ $\leqslant W\left(\mathbb{P},\sum_{k=1}^{N_{C}}p_{k}\hat{\mu}_{k}\right)+W\left(\sum_{k=1}^{N_{C}}p_{k}\hat{\mu}_{k},P(\hat{g}_{3}^{\tau}\circ\hat{g}_{2}\circ\hat{g}_{1}(U_{k}))\right)$
$=W\left(\mathbb{P},\sum_{k=1}^{N_{C}}p_{k}\hat{\mu}_{k}\right)+W\left(P(\hat{g}_{3}^{0}\circ\hat{g}_{2}\circ\hat{g}_{1}(U_{k})),P(\hat{g}_{3}^{\tau}\circ\hat{g}_{2}\circ\hat{g}_{1}(U_{k}))\right)$
@@
where we use (16). ∎
B.3 Proof of results in Section 3.2
@@
Figure 5: Architecture of the deep neural network for $4-$dimensional output. The first (yellow) part approximates the distribution on different connected components using deep ReLu Networks. The remaining two parts are similar to the wide neural network in Figure 1(a).
Before we prove Proposition 1, we need to introduce some important notions. Let $\{\mathcal{C}_{n}\}$ be a sequence of partitions of the unit cube $[0,1]^{d}$. We say $\{\mathcal{C}_{n}\}$ has property (*) if the following properties are satisfied.
@@
Proof of Proposition 1.
Without loss of generality, we can assume that $\mathbb{P}$ is supported on $[0,1]^{d}$, vanishes at the boundary and $\mathbb{P}(B)\geqslant C_{1}\lambda(B)$ for some constant $C_{1}>0$ and all measurable sets $B\in[0,1]^{d}$. Suppose we have proven the conclusion for this case. By 5, there exists $f\in\mathcal{H}([0,1]^{d},K)\cap\mathcal{F}_{L}([0,1]^{d},K)$ such that $f^{-1}_{\#}\mathbb{P}$ satisfies these conditions. There exists continuous function $\gamma$ such that $f^{-1}_{\#}\mathbb{P}=\gamma_{\#}\lambda$, which implies $\mathbb{P}=(f\circ\gamma)_{\#}\lambda$. In particular, if $\gamma$ is $1/d$-Holder continuous, $f\circ\gamma$ is also $1/d$-Holder continuous because $f$ is Lipschitz continuous.
@@
Step 2: Construct Hilbert Curve. By Lemma 1, there exist an sequence of ordering $\{O_{n}\}$ of $\{\mathcal{C}_{n}\}$ that satisfies

### v2

@@ B.1 Illustration of how to construct canonical representations and neural architectures
$\int g(\bm{u},\bm{t},v_{1},\cdots,v_{k})\ d\left(\mu_{V_{k}|\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}}-\widehat{\mu}_{\widehat{V}_{k}|\widehat{\bm{U}}_{0},\bm{T},1:k-1}^{\bm{T}}\right)$
$=g\left(\bm{u},\bm{t},v_{1},\cdots,v_{k-1},f_{k}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)\right)-g\left(\bm{u},\bm{t},v_{1},\cdots,v_{k-1},\widehat{f}_{k}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)\right)$
$\leqslant\left\|f_{k}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)-\widehat{f}_{k}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)\right\|_{\infty}\leqslant S,$ (B.4)
where we have used the Lipschitz property of $g$ in the first inequality. If $V_{k}$ is categorical, let $\widehat{p}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)=(\widehat{p}_{1},\cdots,\widehat{p}_{n_{i}})={\widehat{f}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)}$ and $p\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)=(p_{1},\cdots,p_{n_{i}})=f\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)$ , we get
$\int g(\bm{u},\bm{t},v_{1},\cdots,v_{k})\ d\left(\mu_{V_{k}|\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}}-\widehat{\mu}_{\widehat{V}_{k}|\widehat{\bm{U}}_{0},\bm{T},1:k-1}^{\bm{T}}\right)$
$=\sum_{k=1}^{n_{i}-1}(g(\bm{u},\bm{t},v_{1},\cdots,k)\ -g(\bm{u},\bm{t},v_{1},\cdots,n_{i}))(p_{k}-\widehat{p}_{k})\leqslant K\sum_{k=1}^{n_{i}-1}|p_{k}-\widehat{p}_{k}|\leqslant K^{2}S,$ (B.5)
@@
$=\int g\left(\bm{u},\bm{t},v_{1},\cdots,v_{k-1},{f}_{k}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)\right)\ d\left(\mu_{\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}}-\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:k-1}^{\bm{T}}\right).$ (B.6)
Since $g,{f}_{k}$ are Lipschitz continuous functions, $g\left(\bm{u},\bm{t},v_{1},\cdots,v_{k-1},{f}_{k}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)\right)$ is $(L+1)$-Lipschitz continuous with respect to $(\bm{u},\bm{t},v_{1},\cdots,v_{k-1})$. We have
$\int g\left(\bm{u},\bm{t},v_{1},\cdots,v_{k-1},{f}_{k}\left(\text{pa}(v_{k}),\bm{u}_{V_{k}}\right)\right)$ $d\left(\mu_{\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}}-\widehat{\mu}_{\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}}\right)$
$\leqslant(L+1)W\left(\mu_{\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:k-1}^{\bm{T}}\right).$ (B.7)
@@
$\quad=\sum_{k=1}^{n_{i}}{p}_{k}\int g(\bm{u},\bm{t},x_{1},\cdots,k)\ d\left(\mu_{\bm{U},\bm{T},1:k-1}^{\bm{T}}-\widehat{\mu}_{\widehat{\bm{U}},\bm{T},1:k-1}^{\bm{T}}\right).$
@@
Since $g,f_{k}$ are $L$-Lipschitz continuous functions, $g(\bm{u},\bm{t},v_{1},\cdots,v_{k-1},i)$ is $L$-Lipschitz continuous with respect to $(\bm{u},\bm{t},v_{1},\cdots,v_{k-1})$. We have
$\sum_{k=1}^{n_{i}}\int{p}_{k}g(\bm{u},\bm{t},x_{1},\cdots,k)\ d\left(\mu_{\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}}-\widehat{\mu}_{\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}}\right)$ $\leqslant LW\left(\mu_{\bm{U}_{0},\bm{T},1:k-1}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:k-1}^{\bm{T}}\right).$ (B.8)
Combine (B.4)-(B.8), we prove Equation B.3. By induction, one can easily get
$W\left(\mu_{\bm{U}_{0},\bm{T},1:n_{V}-n_{T}}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:n_{V}-n_{T}}^{\bm{T}}\right)$ $\leqslant(L+1)^{n_{V}-n_{T}}\left(W\left(\mu_{\bm{U}_{0},\bm{T}}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T}}^{\bm{T}}\right)+2K^{2}S/L\right)-2K^{2}S/L$
$=\frac{(L+1)^{n_{V}-n_{T}}-1}{L}2K^{2}S+(L+1)^{n_{V}-n_{T}}W\left(\mu_{\bm{U}_{0},\bm{T}}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T}}^{\bm{T}}\right)$
@@
Let $C_{\mathcal{G}}(L,K)=2K^{2}\cdot\max\{\frac{(L+1)^{n_{V}-n_{T}}-1}{L},(L+1)^{n_{V}-n_{T}}\}$ and notice that $W\left(\mu_{\bm{U}_{0},\bm{T}}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T}}^{\bm{T}}\right)=W\left(P^{\mathcal{M}}(\bm{U}),P^{\widehat{\mathcal{M}}}(\widehat{\bm{U}})\right)$ because interventions $\bm{T}=\bm{t}$ are the same, we get
$W\left(P^{\mathcal{M}}(\bm{V}(\bm{t})),P^{\widehat{\mathcal{M}}}(\bm{V}(\bm{t}))\right)$ $\leqslant W\left(\mu_{\bm{U}_{0},\bm{T},1:n_{V}-n_{T}}^{\bm{T}},\widehat{\mu}_{\widehat{\bm{U}}_{0},\bm{T},1:n_{V}-n_{T}}^{\bm{T}}\right)$
$\leqslant C_{\mathcal{G}}(L,K)\left(S+W\left(P^{\mathcal{M}}(\bm{U}),P^{\widehat{\mathcal{M}}}(\widehat{\bm{U}})\right)\right).$ (B.9)
B.3 Proof of results in Section 3.1
Proof of Theorem 3.5.
@@
Recall that the neural network consists of three parts $\widehat{g}=\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ\widehat{g}_{1}$ and $\widehat{g}_{2},\widehat{g}_{1}$ have a separable form, dealing with each coordinate of the input individually. As mentioned before, each coordinate approximates the distribution of one connected component of the support. We first construct $\widehat{g}_{1}$ and $\widehat{g}_{2}$, then use Gumbel-Softmax layer to combine each component together.
@@
To construct the first two parts, we only need to consider each coordinate individually. For the $i$-th component $C_{i}$, let $\mu_{i}=\mathbb{P}_{C_{i}}/\mathbb{P}(C_{i})$, where $\mathbb{P}_{C_{i}}$ is measure $\mathbb{P}$ restricted to component $C_{i}$. Under 3.4, there exists a Lipschitz map $g_{2}^{i}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$. By [43, Theorem VIII.1], there exists quantized ReLu network $\widehat{g}_{1}^{i}$ of width $W_{1}$ and depth $\Theta(d_{i}^{C})$ (let $s=\frac{1}{\lceil n(n-1)\rceil}$ in [43, Theorem VIII.1]), such that
$W\left(\left((g_{2}^{i})^{-1}\right)_{\#}\mu_{i},P(\widehat{g}_{1}^{i}(U_{i}))\right)\leqslant O\left(W_{1}^{-1/d_{i}^{C}}\right),$
@@
where $U_{i}$ are i.i.d. uniform random variables on $[0,1]$. By [53, Theorem 2], there exist a deep ReLu network $\widehat{g}_{2}^{i,j}$ of width $\Theta\left(d_{i}^{C}\right)$ and depth $L_{2}$ such that
$\|\left(g_{2}^{i}\right)_{j}-\widehat{g}_{2}^{i,j}\|_{\infty}\leqslant O\left(L_{2}^{-2/d_{i}^{C}}\right),$
Let $\widehat{g}_{2}^{i}(x)=\left(\widehat{g}_{2}^{i,1}(x),\cdots,\widehat{g}_{2}^{i,d}(x)\right)$, we get
$\|g_{2}^{i}-\widehat{g}_{2}^{i}\|_{\infty}\leqslant O\left(L_{2}^{-2/d_{i}^{C}}\right).$
Thus, the width of $\widehat{g}_{2}^{i}$ is $\Theta\left(d\cdotp d_{i}^{C}\right)$. Let
$\widehat{g}_{i}(x)=\left(\widehat{g}_{i}^{1}(x_{1}),\cdots,\widehat{g}_{i}^{N_{C}}(x_{N_{C}})\right),i=1,2,x_{k}\in\mathbb{R}^{d_{i}},d_{1}=1,d_{2}=d,k=1,\cdots,N_{C},$
where $N_{C}$ is the number of connected components. By Lemma E.1, we get
$W\left(\mu_{i},P(\widehat{g}_{2}^{i}\circ\widehat{g}_{1}^{i}(U_{i}))\right)$ $=W\left(\left(g_{2}^{i}\right)_{\#}\left(g_{2}^{i}\right)^{-1}_{\#}\mu_{i},\left(\widehat{g}_{2}^{i}\right)_{\#}P(\widehat{g}_{1}^{i}(U_{i}))\right)$
$\leqslant LW\left(\left(g_{2}^{i}\right)^{-1}_{\#}\mu_{i},P(\widehat{g}_{1}^{i}(U_{i}))\right)+\|g_{2}^{i}-\widehat{g}_{2}^{i}\|_{\infty}$
@@
Next, let $p_{i}=\mathbb{P}(C_{i})$ and the distribution of $\widehat{g}_{2}^{k}\circ\widehat{g}_{1}^{k}(U_{k})$ be $\widehat{\mu}_{k}$, then we have $\mathbb{P}=\sum_{k=1}^{N_{C}}p_{k}\mu_{k}$. By Lemma E.2, we have
$W\left(\mathbb{P},\sum_{k=1}^{N_{C}}p_{k}\widehat{\mu}_{k}\right)$ $\leqslant\sum_{k=1}^{N_{C}}p_{k}W\left(\mu_{i},\widehat{\mu}_{k}\right)=O\left(W_{1}^{-1/\max\{d_{i}^{C}\}}+L_{2}^{-2/\max\{d_{i}^{C}\}}\right).$
@@
By Lemma E.3,
$W\left(\left(\nu^{\tau},P(\left(\widehat{g}_{2}^{i}\circ\widehat{g}_{1}^{i}\right)(U_{i}))\right),\left(\nu^{0},P(\left(\widehat{g}_{2}^{i}\circ\widehat{g}_{1}^{i}\right)(U_{i}))\right)\right)\leqslant O\left(\tau-\tau\log\tau\right).$ (B.10)
@@
$\widehat{g}_{3}^{\tau}(x_{1},\cdots,x_{N_{C}})=\sum_{k=1}^{N_{C}}w_{k}^{\tau}\cdotp x_{k},\quad x_{k}\in\mathbb{R}^{d},\ w^{\tau}=\left(w_{1}^{\tau},\cdots,w_{N_{C}}^{\tau}\right).$
We denote $g^{0}_{3}=\lim_{\tau\rightarrow 0}g^{\tau}_{3}$. By 3.4, the support is bounded. Thus, $g_{2}^{i}$ and $\widehat{g}_{2}^{i}$ are bounded, i.e., $\|\widehat{g}_{2}^{i}\ (x)\|_{\infty}\leqslant K$. Since functions $h(w,X)=w^{\mathsf{T}}X$ is Lipschitz continuous in the region $\|w\|_{1}=1,\|X\|_{\infty}\leqslant K$, we get
$W\left(P(\widehat{g}_{3}^{0}\circ\widehat{g}_{2}\circ\widehat{g}_{1}(U_{k})),P(\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ\widehat{g}_{1}(U_{k}))\right)\leqslant O\left(\tau-\tau\log\tau\right)$ (B.11)
from (B.10). Putting things together, we get
$W\left(\mathbb{P},P(\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ\widehat{g}_{1}(U_{k}))\right)$ $\leqslant W\left(\mathbb{P},\sum_{k=1}^{N_{C}}p_{k}\widehat{\mu}_{k}\right)+W\left(\sum_{k=1}^{N_{C}}p_{k}\widehat{\mu}_{k},P(\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ\widehat{g}_{1}(U_{k}))\right)$
$=W\left(\mathbb{P},\sum_{k=1}^{N_{C}}p_{k}\widehat{\mu}_{k}\right)+W\left(P(\widehat{g}_{3}^{0}\circ\widehat{g}_{2}\circ\widehat{g}_{1}(U_{k})),P(\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ\widehat{g}_{1}(U_{k}))\right)$
@@
where we use (B.11). ∎
B.4 Proof of results in Section 3.2
@@
Figure 9: Architecture of the deep neural network for $4-$dimensional output. The first (yellow) part approximates the distribution on different connected components using deep ReLu Networks. The remaining two parts are similar to the wide neural network in Figure 2.
Before we prove Proposition 3.7, we need to introduce some important notions. Let $\{\mathcal{C}_{n}\}$ be a sequence of partitions of the unit cube $[0,1]^{d}$. We say $\{\mathcal{C}_{n}\}$ has property (*) if the following properties are satisfied.
@@
Proof of Proposition 3.7.
Without loss of generality, we can assume that $\mathbb{P}$ is supported on $[0,1]^{d}$, vanishes at the boundary and $\mathbb{P}(B)\geqslant C_{1}\lambda(B)$ for some constant $C_{1}>0$ and all measurable sets $B\in[0,1]^{d}$. Suppose we have proven the conclusion for this case. By 3.6, there exists $f\in\mathcal{H}([0,1]^{d},K)\cap\mathcal{F}_{L}([0,1]^{d},K)$ such that $f^{-1}_{\#}\mathbb{P}$ satisfies these conditions. There exists continuous function $\gamma$ such that $f^{-1}_{\#}\mathbb{P}=\gamma_{\#}\lambda$, which implies $\mathbb{P}=(f\circ\gamma)_{\#}\lambda$. In particular, if $\gamma$ is $1/d$-Holder continuous, $f\circ\gamma$ is also $1/d$-Holder continuous because $f$ is Lipschitz continuous.
@@
Step 2: Construct Hilbert Curve. By Lemma B.2, there exist an sequence of ordering $\{O_{n}\}$ of $\{\mathcal{C}_{n}\}$ that satisfies